OpenStudy (anonymous):

Does this equation has a solution at intervals [-2,10] ?

OpenStudy (anonymous):

\[2^{-x ^{2}}-x=0\]

zepdrix (zepdrix):

So if we add some stuff, then multiply some stuff over, we get, \[\Large\bf 2^{-x^2}\quad=\quad x\qquad\qquad\to\qquad\qquad 1\quad=\quad x\;2^{x^2}\] So maybe it's a little easier to work with from here. We want to know if \(\Large \bf y=1\) intercepts the function \(\Large\bf y=x\;2^{x^2}\) I know we can probably use some calculus techniques to show that if x 2^{x^2} is continuous and is greater than and less than 1 at different points in the interval, then we can probably justify that it has to pass through the line y=1. I'm not sure how we would find the `actual` solution though.. hmm Have you done problems like this before? Is there a special technique? :o

OpenStudy (anonymous):

hm, i do not have the solution, its just says if the a solution can be obtained by using intermediate value theorem. The answer is yes. but how am i suppose to prove that there is indeed a solution without calculating it, if the solution is that complicated.

OpenStudy (anonymous):

how come you cannot apply intermediate value theorem in equations like, 1/x=0, x^2-7x+10=0, x/(x^2-1)=0?

zepdrix (zepdrix):

Well the IVT requires that the function be `continuous` over the given interval. So for the first one: 1/x=0, we would NOT want to apply the IVT to an interval which includes x=0, right?

zepdrix (zepdrix):

Since there is a discontinuity there*

OpenStudy (anonymous):

hmm okay so you're saying, as long as x doesnt equal to 0? so that means the first equation how do u know there is a solution at [-2,10]

zepdrix (zepdrix):

So like if we can show that: f(-2) = (less than 1) `and` f(10) = (greater than 1) `and` the function is continuous `then` the function MUST HAVE passed through y=1 somewhere in the interval.

zepdrix (zepdrix):

So we're talking about something like this. If we can find points like this in our interval,